OpenStudy (anonymous):

I need help on calculus! How do I use related rates? "A spherical balloon is inflated with helium at the rate of 100pi ft^3/min. How fast is the balloon's radius increasing at the instant the radius is 5 ft?"

OpenStudy (lgbasallote):

the formula for volume of a sphere is \[\huge \frac 43 \pi r^3\] right?

OpenStudy (anonymous):

Yup

OpenStudy (lgbasallote):

it says the balloon is inflated with the rate of 100 pi ft^3/min <--this is the rate the volume is increasing in other words, it is the derivaitve of the volume so...take the derivative of 4/3 pi r^3 and equate it to 100 got it?

OpenStudy (anonymous):

ohh thank you so much

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

that will give you the rate of increase of the radius

OpenStudy (anonymous):

but when do i use the 5 ft radius information

OpenStudy (lgbasallote):

so...do you know how to do the rest?

OpenStudy (lgbasallote):

take the derivative of 4/3 pi r^3 first

OpenStudy (anonymous):

so 4r^2pi

OpenStudy (lgbasallote):

close. it's actually \[\huge dV = 4\pi r^2 dr\]

OpenStudy (lgbasallote):

^this is where implicit differentiation is important

OpenStudy (anonymous):

the thing i don't udnerstand is where you got the dV and dr

OpenStudy (lgbasallote):

anyway... dr is the rate the radius is increasing

OpenStudy (lgbasallote):

dV is the derivative of volume

OpenStudy (lgbasallote):

let's try it this way... \[\huge V = \frac 43 \pi r^3\] \[\huge \frac{d}{dr} (V) = 4\pi r^2\] agree?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you multiply each side by dr or something?

OpenStudy (lgbasallote):

now cross multiply... \[\huge dV = 4\pi r^2 dr\]

OpenStudy (anonymous):

yes i understand

OpenStudy (lgbasallote):

yes i did

OpenStudy (lgbasallote):

so anyway... you have \[\huge dV = 4\pi r^2 dr\] you know the value of dV and the value of r...therefore, you can solve for dr

OpenStudy (anonymous):

OHH now i understand thanks :D

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

sorry it took a little long to catch on...i'm only a freshman

OpenStudy (lgbasallote):

no worries. i had difficulties with this one too

OpenStudy (anonymous):

thank you :)

OpenStudy (lgbasallote):

welcome again

OpenStudy (anonymous):

There is an intermediate step here that was sort of ignored. You were never given \(dV\) nor where you supposed to find \(dr\). You were given \(dV/dt\) and asked for \(dr/dt\). So at some point, both sides should have been divided by \(dt\).

OpenStudy (lgbasallote):

i'm really bad with proving that formula.....at least the answer stays teh same though.....